# Platonic solutions of the discrete Nahm equation

### v1

@@ Abstract
The discrete Nahm equation is an integrable nonlinear difference equation for complex $N\times N$ matrices defined on a one-dimensional lattice, with rank and symmetry boundary conditions at the ends of the lattice. Solutions of this system correspond to $SU(2)$ magnetic monopoles of charge $N$ in hyperbolic space, with the curvature related to the number of lattice points. Here some solutions of the discrete Nahm equation are obtained by imposing platonic symmetries, and the spectral curves of the associated hyperbolic monopoles are calculated directly from these solutions.
@@ 1 Introduction
Atiyah [1, 2] observed that for any positive integer $m$, $SU(2)$ magnetic monopoles of charge $N$ in three-dimensional hyperbolic space with sectional curvature $1/m^{2}$ correspond to circle-invariant $SU(2)$ Yang-Mills instantons in four-dimensional Euclidean space with instanton number $mN.$ This correspondence applies in a normalization in which the length of the Higgs field at infinity is equal to $1/2$, but by applying a scaling symmetry this is equivalent to fixing the curvature of hyperbolic space to be 1-1 and setting the length of the Higgs field at infinity equal to $m/2.$ This quantity is often referred to as the monopole mass, hence the hyperbolic monopoles of interest here are known as half-integral mass monopoles.
Braam and Austin [3] adapted the ADHM construction of instantons [4] to half-integral mass hyperbolic monopoles by imposing circle-invariance, to yield a correspondence between these charge $N$ hyperbolic monopoles and an integrable nonlinear difference equation for complex $N\times N$ matrices defined on a one-dimensional lattice with $m+1$ lattice points, together with boundary conditions on the matrices at the ends of the lattice. In the continuum limit, $m\to\infty$, the difference equation becomes the Nahm equation, which appears in adapting the ADHM construction to the flat space limit of monopoles in three-dimensional Euclidean space [5]. As the Braam-Austin difference equation is an integrable discretization of the Nahm equation it is commonly referred to as the discrete Nahm equation.
@@
In this paper the first examples of solutions of the discrete Nahm equation with $m>1$ and $N>2$ are presented, by applying platonic symmetries to simplify the system. The associated monopole spectral curves are obtained directly from these solutions. Some, but not all, of these spectral curves have been obtained previously using methods of algebraic geometry [12]. Section 2 provides a brief review of the discrete Nahm equation. Section 3 describes the imposition of platonic symmetry and presents the lowest charge examples with tetrahedral, octahedral and icosahedral symmetry, with $N=3,4,7,$ respectively.
@@ 2 The discrete Nahm equation
$B_{2\ell+2}=W_{2\ell+1}^{-1}B_{2\ell}W_{2\ell+1},\qquad\quad\ \ \qquad$ $\mbox{ for }\ \ell=0,1,..(m-3)/2,$ (2.1)
$W_{2\ell+1}W_{2\ell+1}^{\dagger}=W_{2\ell-1}^{\dagger}W_{2\ell-1}+[B_{2\ell}^{\dagger},B_{2\ell}],\qquad$ $\mbox{ for }\ \ell=0,1,..(m-1)/2,$ (2.2)
@@ 3 Platonic solutions
where (3.1) and (3.5) have been used to obtain this expression by exchanging the $\alpha$ and $\beta$ dependence inside the determinant for conjugation of the $Y_{i}$ by $F_{\cal O}$. This proves that the curve is $G$-symmetric. Finally, the parameters in the $G$-symmetric triplet $(Y_{1},Y_{2},Y_{3})$ are determined by evolving the initial data, (3.2) and (3.3), to obtain $W_{m}$ and imposing the condition that this has rank one. This provides the required platonic solution of the discrete Nahm equation and the associated spectral curve. The remainder of this section provides some examples of applying this procedure.
@@ 3.1 $N=3$ tetrahedral solutions
$m$ $c_{T}$ $c_{O}$ $c_{Y}$
1 0.5774 0.3333 2.0000
3 0.1475 0.0718 0.3333
5 0.0593 0.0243 0.0902
7 0.0297 0.0105 0.0317
9 0.0170 0.0053 0.0132
11 0.0106 0.0029 0.0062
13 0.0071 0.0018 0.0032
Table 1: The spectral curve coefficients, $c_{T},c_{O},c_{Y}$, to 4 decimal places, for the odd values of $m$ from 1 to 13.

### v2

@@ Abstract
The discrete Nahm equation is an integrable nonlinear difference equation for complex $N\times N$ matrices defined on a one-dimensional lattice, with rank and symmetry boundary conditions at the ends of the lattice. Solutions of this system correspond to $SU(2)$ magnetic monopoles of charge NN in hyperbolic space, with the curvature related to the number of lattice points. Here some solutions of the discrete Nahm equation are obtained by imposing platonic symmetries, and the spectral curves of the associated hyperbolic monopoles are calculated directly from these solutions.
@@ 1 Introduction
Atiyah [1, 2] observed that for any positive integer $m$, $SU(2)$ magnetic monopoles of charge NN in three-dimensional hyperbolic space with sectional curvature $1/m^{2}$ correspond to circle-invariant $SU(2)$ Yang-Mills instantons in four-dimensional Euclidean space with instanton number $mN.$ This correspondence applies in a normalization in which the length of the Higgs field at infinity is equal to $1/2$, but by applying a scaling symmetry this is equivalent to fixing the curvature of hyperbolic space to be 1-1 and setting the length of the Higgs field at infinity equal to $m/2.$ This quantity is often referred to as the monopole mass, hence the hyperbolic monopoles of interest here are known as half-integral mass monopoles.
Braam and Austin [3] adapted the ADHM construction of instantons [4] to half-integral mass hyperbolic monopoles by imposing circle-invariance, to yield a correspondence between these charge NN hyperbolic monopoles and an integrable nonlinear difference equation for complex $N\times N$ matrices defined on a one-dimensional lattice with $m+1$ lattice points, together with boundary conditions on the matrices at the ends of the lattice. In the continuum limit, $m\to\infty$, the difference equation becomes the Nahm equation, which appears in adapting the ADHM construction to the flat space limit of monopoles in three-dimensional Euclidean space [5]. As the Braam-Austin difference equation is an integrable discretization of the Nahm equation it is commonly referred to as the discrete Nahm equation.
@@
In this paper the first examples of solutions of the discrete Nahm equation with $m>1$ and $N>2$ are presented, by applying platonic symmetries to simplify the system. The associated monopole spectral curves are obtained directly from these solutions. Some, but not all, of these spectral curves have been obtained previously using methods of algebraic geometry [12]. Section 2 provides a brief review of the discrete Nahm equation. Section 3 describes the imposition of platonic symmetry and presents some examples, including the lowest charge solutions with tetrahedral, octahedral and icosahedral symmetry, with $N=3,4,7,$ respectively.
@@ 2 The discrete Nahm equation
$B_{2\ell+2}=W_{2\ell+1}^{-1}B_{2\ell}W_{2\ell+1},\qquad\quad\ \ \qquad$ $\mbox{ for }\ \ell=0,1,..,(m-3)/2,$ (2.1)
$W_{2\ell+1}W_{2\ell+1}^{\dagger}=W_{2\ell-1}^{\dagger}W_{2\ell-1}+[B_{2\ell}^{\dagger},B_{2\ell}],\qquad$ $\mbox{ for }\ \ell=0,1,..,(m-1)/2,$ (2.2)
@@ 3 Platonic solutions
where (3.1) and (3.5) have been used to obtain this expression by exchanging the $\alpha$ and $\beta$ dependence inside the determinant for conjugation of the $Y_{i}$ by $F_{\cal O}$. This proves that the curve is $G$-symmetric. Finally, the parameters in the $G$-symmetric triplet $(Y_{1},Y_{2},Y_{3})$ are determined by evolving the initial data, (3.2) and (3.3), to obtain $W_{m}$ and imposing the condition that this has rank one. This provides the required platonic solution of the discrete Nahm equation and the associated spectral curve.
The remainder of this section provides some examples of applying this procedure. As mentioned earlier, the lowest charge solutions with platonic symmetry groups $G=T,O,Y$ are $N=3,4,7.$ This can be established using a correspondence between hyperbolic monopoles and rational maps between Riemann spheres [17], to convert the question of the existence of a $G$-symmetric hyperbolic monopole with charge NN to that of a $G$-symmetric rational map of degree NN. This can be determined using a little representation theory [18], as follows. Let $\underline{N+1}$ denote the $(N+1)$-dimensional irreducible representation of $SU(2)$, and $\underline{N+1}|_{G^{*}}$ its restriction to the binary platonic subgroup $G^{*}\subset SU(2).$ In general $\underline{N+1}|_{G^{*}}$ will be a reducible representation of $G^{*}$, and the existence of a symmetric rational map requires that it contains a 2-dimensional irreducible representation of $G^{*}$ or at least two 1-dimensional irreducible representations. Applying this criterion reveals the above charges as the lowest possible for the given platonic symmetries. The same analysis confirms the existence of a charge $N=5$ hyperbolic monopole with octahedral symmetry. Solutions of the discrete Nahm equation corresponding to all four of these platonic monopoles, with charges $N=3,4,5,7,$ will be presented below. The representation analysis shows that there are no platonic solutions with charge $N=6$.
@@ 3.1 $N=3$ tetrahedral solutions
$m$ $c_{T}$ $c_{O}$ $\widetilde{c}_{O}$ $c_{Y}$
1 0.5774 0.3333 1.0000 2.0000
3 0.1475 0.0718 0.2000 0.3333
5 0.0593 0.0243 0.0704 0.0902
7 0.0297 0.0105 0.0316 0.0317
9 0.0170 0.0053 0.0164 0.0132
11 0.0106 0.0029 0.0094 0.0062
13 0.0071 0.0018 0.0057 0.0032
Table 1: The spectral curve coefficients, $c_{T},c_{O},\widetilde{c}_{O},c_{Y}$, to 4 decimal places, for the odd values of $m$ from 1 to 13.
@@ 3.2 $N=4$ octahedral solutions
3.3 $N=5$ octahedral solutions
A symmetric triplet with $N=5$ and $G=O$ is given by
$Y_{1}+iY_{2}=\frac{d}{4}\left(\begin{array}[]{ccccc}0&\sqrt{2}&0&-i\sqrt{2}&0\\ \sqrt{2}&0&2\sqrt{3}&0&\sqrt{2}\\ 0&2\sqrt{3}&0&i2\sqrt{3}&0\\ -i\sqrt{2}&0&i2\sqrt{3}&0&-i\sqrt{2}\\ 0&\sqrt{2}&0&-i\sqrt{2}&0\end{array}\right),\qquad Y_{3}=d\left(\begin{array}[]{ccccc}-1&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&1\end{array}\right),$ (3.27)
with $d\in(0,1)$. Applying (3.2) and (3.3), this triplet generates the matrices
$B_{0}=\frac{d}{4}\left(\begin{array}[]{ccccc}0&\frac{\sqrt{2}}{\sqrt{1+d}}&0&\frac{-i\sqrt{2}}{\sqrt{1+d}}&0\\ \frac{\sqrt{2}}{\sqrt{1+d}}&0&2\sqrt{3}&0&\frac{\sqrt{2}}{\sqrt{1-d}}\\ 0&2\sqrt{3}&0&i2\sqrt{3}&0\\ \frac{-i\sqrt{2}}{\sqrt{1+d}}&0&i2\sqrt{3}&0&\frac{-i\sqrt{2}}{\sqrt{1-d}}\\ 0&\frac{\sqrt{2}}{\sqrt{1-d}}&0&\frac{-i\sqrt{2}}{\sqrt{1-d}}&0\end{array}\right),$ (3.28)
[…]
where $\widetilde{q}=4-d^{2}-4d.$ Defining
$\widetilde{c}_{O}=\frac{3d^{4}}{4(1-d^{2})}$ (3.30)
the octahedrally symmetric spectral curve is
$(\eta-\zeta)^{5}-\widetilde{c}_{O}(\eta-\zeta)(\eta^{4}\zeta^{4}+6\eta^{2}\zeta^{2}+4\eta\zeta(\eta^{2}+\zeta^{2})+1)=0,$ (3.31)
invariant under the generators (3.22) of the octahedral group. The determinant of (3.29) is
$\mbox{det}(W_{1})=\frac{(2-3d^{2})^{2}}{4(1-d^{2})},$ (3.32)
with $W_{1}$ having rank one if $d$ is equal to $d_{1}=\sqrt{\frac{2}{3}}$, so that $\widetilde{c}_{O}=1.$ This provides the $m=1$ solution of the discrete Nahm equation, with a spectral curve in agreement with that obtained using the JNR approach [8] that is applicable for $m=1$. If $d<d_{1}$ the system is evolved along the lattice to produce matrices $B_{2}$ and $W_{3}$ with
$\mbox{det}(W_{3})=\frac{(2+d^{2})(2-5d^{2})^{2}}{4(2-3d^{2})(1-d^{2})},$ (3.33)
which vanishes for $d$ equal to $d_{3}=\sqrt{\frac{2}{5}}$. The matrices for this $m=3$ solution are
$B_{2}=\left(\begin{array}[]{ccccc}0&\frac{\sqrt{15}}{10\sqrt{9-2\sqrt{10}}}&0&\frac{-1}{2\sqrt{9-2\sqrt{10}}}i&0\\ \frac{\sqrt{15}\left(3-\sqrt{10}\right)}{6\sqrt{9-2\sqrt{10}}}&0&\frac{3\sqrt{10}}{20}&0&\frac{5\sqrt{6}}{\sqrt{9-2\sqrt{10}}\,\left(20-4\sqrt{10}\right)}\\ 0&\frac{\sqrt{10}}{4}&0&\frac{\sqrt{6}}{4}i&0\\ \frac{\left(-3+\sqrt{10}\right)}{2\sqrt{9-2\sqrt{10}}}i&0&\frac{\sqrt{6}}{4}i&0&\frac{3\sqrt{10}}{\sqrt{9-2\sqrt{10}}\,\left(-20+4\sqrt{10}\right)}i\\ 0&\frac{\sqrt{6}\,\left(5-\sqrt{10}\right)}{20\sqrt{9-2\sqrt{10}}}&0&\frac{\left(-5+\sqrt{10}\right)\sqrt{10}}{20\sqrt{9-2\sqrt{10}}}i&0\end{array}\right),$ (3.34)
and the rank one matrix
$W_{3}=\frac{\sqrt{9-2\sqrt{10}}}{205}\left(\begin{array}[]{ccccc}85-13\sqrt{10}&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\\ -150-47\sqrt{10}&0&0&0&0\end{array}\right).$ (3.35)
Substituting $d=d_{3}$ into (3.30) gives the $m=3$ spectral curve coefficient $\widetilde{c}_{O}=\frac{1}{5}$. Continuing the evolution along the lattice for $d<d_{3}$, and calculating the values of $d$ with $\mbox{det}(W_{m})=0,$ produces the numerical values of $\widetilde{c}_{O}$ presented in Table 1 for larger values of $m$.